OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 3x + 2y = 6 and y = -3/2 x - 1 A. (2, 0) B. (1, 1.5) C. no solution D. infinitely many solutions

hero (hero):

y = (6 - 3x)/2 y = -3x/2 - 1 y = y 2(-3x/2 - 1) = 6 - 3x -6x/2 -2 = 6 - 3x -6x/2 = 8 - 3x -6x = 16 - 6x No Solution

OpenStudy (anonymous):

thanks :) <3 it helped

hero (hero):

Good